OpenStudy (anonymous):

Which of these is a simplified form of the equation 7y + 8 = 9 + 3y + 2y? 7y = 6 2y = 1 12y = 17 5y = 17

OpenStudy (anonymous):

Hmm,what do u think? @loljanis

OpenStudy (anonymous):

I think it would be A.

OpenStudy (is3535):

implifying 7y + 8 = 9 + 3y + 2y Reorder the terms: 8 + 7y = 9 + 3y + 2y Combine like terms: 3y + 2y = 5y 8 + 7y = 9 + 5y Solving 8 + 7y = 9 + 5y Solving for variable 'y'. Move all terms containing y to the left, all other terms to the right. Add '-5y' to each side of the equation. 8 + 7y + -5y = 9 + 5y + -5y Combine like terms: 7y + -5y = 2y 8 + 2y = 9 + 5y + -5y Combine like terms: 5y + -5y = 0 8 + 2y = 9 + 0 8 + 2y = 9 Add '-8' to each side of the equation. 8 + -8 + 2y = 9 + -8 Combine like terms: 8 + -8 = 0 0 + 2y = 9 + -8 2y = 9 + -8 Combine like terms: 9 + -8 = 1 2y = 1 Divide each side by '2'. y = 0.5 Simplifying y = 0.5

OpenStudy (is3535):

ill hope i help

OpenStudy (anonymous):

oh okay, thank you so much! it really did help!

OpenStudy (study_together):

7y + 8 = 9 + 3y + 2y? It is C because if we simplify the equation after the equal sign we get 9+5y so then we add the 7y and the 5y together to get 12y and also we add the 9 and the 8 together to get 17